On October 1, 2018, Renfro Company purchased to hold maturity, 4,000, $1,000, 9% Bonds for $3,900,000. The bonds, which mature on February 1, 2027, pay interest semiannually on February 1 and August 1. Renfro uses the straight-line method of amortization. The fair value of the bonds at December 31, 2018 was $3,960,000. Renfro chooses the fair value option. The bonds should be reported in the December 31, 2018 balance sheet at
Answer #1

The value of the bonds ( investment ) = 3900,000 - 3900,000 *9%*2/12= 38,41,500 ( since $58500 is the accrued interest , it is not the value of bonds )

The bonds are sold at a discount of 4000000 - 38,41,500 = $158500

Now, the period of the bonds is October 1, 2018 to February 1, 2027 i.e 100 months

So, discount amortized from October 1, 2018 to December 31 , 2018 = 158500 * 3 / 100 = $4755

The carrying value of the bonds on December 31 , 2018 = 3841500 + 4755 = $3846255